A water tank is formed with the combination of cylinder and hemisphere the total height of the tank is 24 metre and a base area is 616 sq. m. if the tank is filled with water at the rate of 65 paise per litre what is the total cost of water find it

Grade:10

2 Answers

Arun
4 years ago
Base area = 616m sq.
πr^2 = 616
r^2 = 616*7/22
r^2 = 196
r = 14
Radius =14 m
Height = 48 m
CSA of cylinder = 2πrh
=2*22/7*14*48
=5984m sq.

Volume of cylinder = πr^2h
=22/7*14*14*48
=29568m cube
Vikas TU
4 years ago
Dear student 
Let me complete the question , ans ans is incomplete.
29568 cubic metre = 29568000 Litre
1 litre cost = 65 paisa = 0.65 Rupees
so 29568000 Litre = 1,92,77,700 Rupees 
Hope this helps 
Good Luck